How do I approach this?*58. a) If g is a given function which is continuous and positive on the interval (0, L], show that the only solution of the
Answer #1

yto 2 ->1,17 (b) y+2y taxy yoosa yil) 20 Arume that then find ? ytry texy=o (D?+20+2x)g=0 → yox) = Gear+Gedix Il Mar - 2+(1) #C) Y - gay =0(1), y(0)= y(L) 20 s txalol Suppose yo on (0,L3Since y is cleffevenhalle funchon st (i) y el contenuvas un